OpenStudy (anonymous):

If 15.0g of Co2 gas has a volume of 3.0L at 300K, what is its pressure in mmHg?

OpenStudy (anonymous):

First, convert 15.0g of \(\rm CO_2\) to moles. \(\rm {n=\large{\frac{15.0}{44.01}} \approx \normalsize 0.341 ~moles}\) Then apply \(\rm n=\large \frac{PV}{RT}~\rightarrow \normalsize P\large= \frac{nRT}{V}=\frac{0.341\times 0.0821 \times 300}{3.0}\approx \normalsize 2.8~atm\) Now that you need the pressure in mmHg, convert it by using \(\rm P=2.8~atm\times \large \frac{760~mmHg}{1~atm} \normalsize \approx 2100 ~mmHg\) Hope it helps

OpenStudy (anonymous):

Do you get everything?

OpenStudy (anonymous):

yes thank you!

OpenStudy (anonymous):

Great

OpenStudy (anonymous):

And you're welcome